OpenStudy (anonymous):

Which of the following possibilities will form a triangle? Side = 15 cm, side = 6 cm, side = 8 cm Side = 15 cm, side = 6 cm, side = 9 cm Side = 16 cm, side = 9 cm, side = 6 cm Side = 16 cm, side = 9 cm, side = 8 cm

OpenStudy (anonymous):

@ChillOut @campbell_st

OpenStudy (anonymous):

@triciaal

OpenStudy (compassionate):

The two smaller sides need to make up the larger side when squared. The formula x^2 + b^2 = c^2 shows this. Plug in your known values and see if its true. If a and b = c, then yes, it forms a triangle.

OpenStudy (anonymous):

For triangles, the shortest 2 sides must add up and be longer than the 3rd side. For each option add the shortest two sides and compare it to the longest side.

OpenStudy (anonymous):

ok then hold on ..

OpenStudy (anonymous):

would it be D.?

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

9^2 + 8^2 = 16^2 145 = 256 No. That's not it.

OpenStudy (anonymous):

what is ^ lol

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

its A right? @Compassionate

OpenStudy (anonymous):

nvm i mean B not A

OpenStudy (compassionate):

a + b > c So your two sides, added together, must be greater than c.

OpenStudy (compassionate):

Side = 16 cm, side = 9 cm, side = 8 cm 16 + 9 > 8 9 + 8 > 16 8 + 16 > 9 Therefore, your answer comes out to D. Understand?

OpenStudy (compassionate):

I was trying to poke around and see if you'd understand. Basically, if all points given are greater than c, then it can form a triangle.

OpenStudy (compassionate):

I wanted to see if you were paying attention or just going along with what I was saying

OpenStudy (anonymous):

oh ok

OpenStudy (xcharlizea):

add all three niumbers and see if any adds up to 180 then that woud be your answer

OpenStudy (anonymous):

well then thank you!!!

OpenStudy (compassionate):

@xcharlizea , that's wrong. The formula is, you take the three points, and if a + b > c, then it forms a triangle, but you have to do this for every possible arrangement. So, if you had the points, 1cm, 2cm, 3cm 1cm + 2cm > 3cm (true) 2cm + 3cm > 1cm (true) 3cm + 1cm > 2cm (true) So the points 1cm, 2cm, 3cm will form a triangle.

OpenStudy (xcharlizea):

@Compassionate oo right thats what i meant lol